OpenStudy (anonymous):

HELP ME PLEASE , ILL MEDAL Which equations in point-slope form are equations of the line that pass through the points (4, 5) and (−3, −1)? Choose exactly two answers that are correct. A. y - 3 = 7/6 ( x - 1 ) B. y -5 = 6/7 ( x -4 ) C. y + 1 = 6/7 ( x + 3 ) D. y - 4 = 7/6 ( x - 5 )

OpenStudy (anonymous):

@ccswims

OpenStudy (ccswims):

first your have to find the slope

OpenStudy (ccswims):

so do your average rate of change formula: -1 - 5/-3 - 4

OpenStudy (ccswims):

You get 6/7... Which only leaves answer B and C left. Point-slope form = y - y1 = m (x - x1)

OpenStudy (ccswims):

I mean B!

OpenStudy (ccswims):

B is the right answer :)

OpenStudy (anonymous):

thats the only right answer ?? cause it say you can pick all the is correct @ccswims

OpenStudy (anonymous):

you still on ?? @ccswims